Let x and y be real numbers satisfying the eqn x2 - 4x + y2 + 3 =0. If the max and min values of x2+y2 are a and b respectively, then show that the numerical value of a-b is 8

Grade:12

1 Answers

Ramesh V
14 years ago

the given equation is a circle with centre (2,0) and radius of 1 unit

(x-2)2+x2=1

the genral point on circle in polar form is : ( 2+cos x , sin x )

on putting in x2+y2 ,

(2+cos x)2+(sin x)2 = -4.sin x

sinx has max. and min. values at 1 , -1

-4 < (2+cos x)2+(sin x)< 4

the numerical value: a-b = 8

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh  
IIT Kgp - 05 batch